BRICS Dissertation Series, 1996–05

January 16, 2007

This document is also available as PostScript, DVI,

References

[DS-05-9] Henning Korsholm Rohde. Formal Aspects of Partial Evaluation. December 2005. PhD thesis. xii+188.

[DS-05-8] Dariusz Biernacki. The Theory and Practice of Programming Languages with Delimited Continuations. December 2005. PhD thesis.

[DS-05-7] Kasper Dupont. Disk Encryption, Group Identification, Byzantine Agreement, and Threshold RSA. December 2005. PhD thesis.

[DS-05-6] Saurabh Agarwal. GCD Algorithms for Quadratic Number Rings. October 2005. PhD thesis.

[DS-05-5] Jesus Fernando Almansa Guerra. A Study for Cryptologic Protocols. September 2005. PhD thesis.

[DS-05-4] Bolette Ammitzbøll Madsen. Exact Algorithms and Exact Satisfiability. September 2005. PhD thesis.

[DS-05-3] Marco Carbone. Trust and Mobility. 2005. PhD thesis.

[DS-05-2] Jørgen Iversen. Formalisms and tools supporting Constructive Action Semantics. May 2005. PhD thesis. xii+189 pp.

[DS-05-1] Abstract, PostScript, PDF.

Kirill Morozov. On Cryptographic Primitives Based on Noisy Channels. March 2005. PhD thesis. xii+102 pp.

[DS-04-6] Abstract, PostScript, PDF.

Paweł Sobocinski. Deriving Process Congruences from Reaction Rules. December 2004. PhD thesis. xii+216 pp.

[DS-04-5] Abstract, PostScript, PDF.

Bjarke Skjernaa. Exact Algorithms for Variants of Satisfiability and Colouring Problems. November 2004. PhD thesis. x+112 pp.

[DS-04-4] Jesper Makholm Byskov. Exact Algorithms for Graph Colouring and Exact Satisfiability. November 2004. PhD thesis.

[DS-04-3] Abstract, PostScript, PDF.

Jens Groth. Honest Verifier Zero-knowledge Arguments Applied. October 2004. PhD thesis. xii+119 pp.

[DS-04-2] Alex Rune Berg. Rigidity of Frameworks and Connectivity of Graphs. July 2004. PhD thesis. xii+173 pp.

[DS-04-1] Abstract, PostScript, PDF.

Bartosz Klin. An Abstract Coalgebraic Approach to Process Equivalence for Well-Behaved Operational Semantics. May 2004. PhD thesis. x+152 pp.

[DS-03-14] Abstract, PostScript, PDF.

Daniele Varacca. Probability, Nondeterminism and Concurrency: Two Denotational Models for Probabilistic Computation. November 2003. PhD thesis. xii+163 pp.

[DS-03-13] Abstract, PostScript, PDF.

Mikkel Nygaard. Domain Theory for Concurrency. November 2003. PhD thesis. xiii+161 pp.

[DS-03-12] Abstract, PostScript, PDF.

Paulo B. Oliva. Proof Mining in Subsystems of Analysis. September 2003. PhD thesis. xii+198 pp.

[DS-03-11] Maciej Koprowski. Cryptographic Protocols Based on Root Extracting. August 2003. PhD thesis. xii+138 pp.

[DS-03-10] Abstract, PostScript, PDF.

Serge Fehr. Secure Multi-Player Protocols: Fundamentals, Generality, and Efficiency. August 2003. PhD thesis. xii+125 pp.

[DS-03-9] Abstract, PostScript, PDF.

Mads J. Jurik. Extensions to the Paillier Cryptosystem with Applications to Cryptological Protocols. August 2003. PhD thesis. xii+117 pp.

[DS-03-8] Abstract, PostScript, PDF.

Jesper Buus Nielsen. On Protocol Security in the Cryptographic Model. August 2003. PhD thesis. xiv+341 pp.

[DS-03-7] Abstract, PostScript, PDF.

Mario José Cáccamo. A Formal Calculus for Categories. June 2003. PhD thesis. xiv+151.

[DS-03-6] Abstract, PostScript, PDF.

Rasmus K. Ursem. Models for Evolutionary Algorithms and Their Applications in System Identification and Control Optimization. June 2003. PhD thesis. xiv+183 pp.

[DS-03-5] Giuseppe Milicia. Applying Formal Methods to Programming Language Design and Implementation. June 2003. PhD thesis. xvi+211.

[DS-03-4] Abstract, PostScript, PDF.

Federico Crazzolara. Language, Semantics, and Methods for Security Protocols. May 2003. PhD thesis. xii+160.

[DS-03-3] Abstract, PostScript, PDF.

Jiří Srba. Decidability and Complexity Issues for Infinite-State Processes. 2003. PhD thesis. xii+171 pp.

[DS-03-2] Frank D. Valencia. Temporal Concurrent Constraint Programming. February 2003. PhD thesis. xvii+174.

[DS-03-1] Abstract, PostScript, PDF.

Claus Brabrand. Domain Specific Languages for Interactive Web Services. January 2003. PhD thesis. xiv+214 pp.

[DS-02-5] Abstract, PostScript, PDF.

Rasmus Pagh. Hashing, Randomness and Dictionaries. October 2002. PhD thesis. x+167 pp.

[DS-02-4] Abstract, PostScript, PDF.

Anders Møller. Program Verification with Monadic Second-Order Logic & Languages for Web Service Development. September 2002. PhD thesis. xvi+337 pp.

[DS-02-3] Abstract, PostScript, PDF.

Riko Jacob. Dynamic Planar Convex hull. May 2002. PhD thesis. xiv+112 pp.

[DS-02-2] Abstract, PostScript, PDF.

Stefan Dantchev. On Resolution Complexity of Matching Principles. May 2002. PhD thesis. xii+68 pp.

[DS-02-1] Abstract, PDF.

M. Oliver Möller. Structure and Hierarchy in Real-Time Systems. April 2002. PhD thesis. xvi+228 pp.

[DS-01-10] Abstract, PostScript, PDF.

Mikkel T. Jensen. Robust and Flexible Scheduling with Evolutionary Computation. November 2001. PhD thesis. xii+299 pp.

[DS-01-9] Abstract, PostScript, PDF.

Flemming Friche Rodler. Compression with Fast Random Access. November 2001. PhD thesis. xiv+124 pp.

[DS-01-8] Niels Damgaard. Using Theory to Make Better Tools. October 2001. PhD thesis.

[DS-01-7] Lasse R. Nielsen. A Study of Defunctionalization and Continuation-Passing Style. August 2001. PhD thesis. iv+280 pp.

[DS-01-6] Abstract, PostScript, PDF.

Bernd Grobauer. Topics in Semantics-based Program Manipulation. August 2001. PhD thesis. ii+x+186 pp.

[DS-01-5] Abstract, PostScript, PDF.

Daniel Damian. On Static and Dynamic Control-Flow Information in Program Analysis and Transformation. August 2001. PhD thesis. xii+111 pp.

[DS-01-4] Abstract, PostScript, PDF.

Morten Rhiger. Higher-Order Program Generation. August 2001. PhD thesis. xiv+144 pp.

[DS-01-3] Abstract, PostScript, PDF.

Thomas S. Hune. Analyzing Real-Time Systems: Theory and Tools. March 2001. PhD thesis. xii+265 pp.

[DS-01-2] Jakob Pagter. Time-Space Trade-Offs. March 2001. PhD thesis. xii+83 pp.

[DS-01-1] Abstract, PostScript, PDF.

Stefan Dziembowski. Multiparty Computations — Information-Theoretically Secure Against an Adaptive Adversary. January 2001. PhD thesis. 109 pp.

[DS-00-7] Abstract, PostScript, PDF.

Marcin Jurdziński. Games for Verification: Algorithmic Issues. December 2000. PhD thesis. ii+112 pp.

[DS-00-6] Abstract, PostScript, PDF.

Jesper G. Henriksen. Logics and Automata for Verification: Expressiveness and Decidability Issues. May 2000. PhD thesis. xiv+229 pp.

[DS-00-5] Abstract, PostScript, PDF.

Rune B. Lyngsø. Computational Biology. March 2000. PhD thesis. xii+173 pp.

[DS-00-4] Abstract, PostScript, PDF.

Christian N. S. Pedersen. Algorithms in Computational Biology. March 2000. PhD thesis. xii+210 pp.

[DS-00-3] Theis Rauhe. Complexity of Data Structures (Unrevised). March 2000. PhD thesis. xii+115 pp.

[DS-00-2] Abstract, PostScript, PDF.

Anders B. Sandholm. Programming Languages: Design, Analysis, and Semantics. February 2000. PhD thesis. xiv+233 pp.

[DS-00-1] Abstract, PostScript, PDF.

Thomas Troels Hildebrandt. Categorical Models for Concurrency: Independence, Fairness and Dataflow. February 2000. PhD thesis. x+141 pp.

[DS-99-1] Abstract, PostScript, PDF.

Gian Luca Cattani. Presheaf Models for Concurrency (Unrevised). April 1999. PhD thesis. xiv+255 pp.

[DS-98-3] Abstract, PostScript, PDF. Kim Sunesen. Reasoning about Reactive Systems. December 1998. PhD thesis. xvi+204 pp.

[DS-98-2] Abstract, PostScript, PDF. Søren B. Lassen. Relational Reasoning about Functions and Nondeterminism. December 1998. PhD thesis. x+126 pp.

[DS-98-1] Abstract, PostScript, PDF. Ole I. Hougaard. The CLP(OIH) Language. February 1998. PhD thesis. xii+187 pp.

[DS-97-3] Abstract, PostScript, PDF. Thore Husfeldt. Dynamic Computation. December 1997. PhD thesis. 90 pp.

[DS-97-2] Abstract, PostScript, PDF. Peter Ørbæk. Trust and Dependence Analysis. July 1997. PhD thesis. x+175 pp.

[DS-97-1] Abstract, PostScript, PDF. Gerth Stølting Brodal. Worst Case Efficient Data Structures. January 1997. PhD thesis. x+121 pp.

[DS-96-4] Abstract, PostScript, PDF. Torben Braüner. An Axiomatic Approach to Adequacy. November 1996. Ph.D. thesis. 168 pp.

[DS-96-3] Abstract, PostScript, PDF. Lars Arge. Efficient External-Memory Data Structures and Applications. August 1996. Ph.D. thesis. xii+169 pp.

[DS-96-2] Abstract, PostScript, PDF. Allan Cheng. Reasoning About Concurrent Computational Systems. August 1996. Ph.D. thesis. xiv+229 pp.

[DS-96-1] Abstract, PostScript, PDF. Urban Engberg. Reasoning in the Temporal Logic of Actions — The design and implementation of an interactive computer system. August 1996. Ph.D. thesis. xvi+222 pp.